# Privileged Information Distillation for Language Models

### v1

@@ Abstract
Training-time privileged information (PI) can enable language models to succeed on tasks they would otherwise fail, making it a powerful tool for reinforcement learning in hard, long-horizon settings. However, transferring capabilities learned with PI to policies that must act without it at inference time remains a fundamental challenge. We study this problem in the context of distilling frontier models for multi-turn agentic environments, where closed-source systems typically hide their internal reasoning and expose only action trajectories. This breaks standard distillation pipelines, since successful behavior is observable but the reasoning process is not. For this, we introduce $\pi$-Distill, a joint teacher-student objective that trains a PI-conditioned teacher and an unconditioned student simultaneously using the same model. Additionally, we also introduce On-Policy Self-Distillation (OPSD), an alternative approach that trains using Reinforcement Learning (RL) with a reverse KL-penalty between the student and the PI-conditioned teacher. We show that both of these algorithms effectively distill frontier agents using action-only PI. Specifically we find that $\pi$-Distill and in some cases OPSD, outperform industry standard practices (Supervised finetuning followed by RL) that assume access to full Chain-of-Thought supervision across multiple agentic benchmarks, models, and forms of PI. We complement our results with extensive analysis that characterizes the factors enabling effective learning with PI, focusing primarily on $\pi$-Distill and characterizing when OPSD is competitive.
Figure 1: Overview of the $\pi$-Distill framework. (1) Successful trajectories (not shown) are collected from a frontier agent that exposes only actions while hiding its Chain-of-Thought. (2) These trajectories are transformed into training-time privileged information (PI) and used to sample a PI-conditioned teacher policy $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})$}. (3) The PI-conditioned teacher and an unconditioned student $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}(\mathbf{o}|s)$} share parameters and are trained jointly, enabling transfer of privileged knowledge to a test-time policy that acts without PI.
@@ 1 Introduction
Training-time PI can be particularly useful for Reinforcement Learning (RL) with LMs, where learning is contingent on the model’s ability to first succeed on a task (zelikman2022starbootstrappingreasoningreasoning; zelikman2024quietstarlanguagemodelsteach). That is, the model can leverage train-time PI to succeed at tasks it would otherwise fail, effectively enabling it to bootstrap its learning from these successful experiences. As noted above, the underlying challenge is transfer: where training a model with PI and obtaining a test-time policy that retains those enhanced capabilities without PI.
@@
We ground our work in the task of distilling frontier models for complex multi-turn agentic settings. Typically, the industry standard for these tasks involves Supervised Fine-Tuning (SFT) on frontier model outputs followed by Reinforcement Learning (RL). Unfortunately, it has become common practice for closed-source frontier models to hide their full Chain-of-Thought (CoT) reasoning, at best producing a summary alongside the action they intend to take (openai2024openaio1card). This opacity undermines standard distillation methods, as we can observe what successful agents do but not their reasoning.
@@
Contributions 1. Algorithms for Privileged Information: We introduce $\pi$-Distill and On-Policy Self-Distillation (OPSD), two novel methods for training policies using train-time privileged information. 2. Distilling without CoT: We show that our methods effectively distill frontier models using actions alone, bypassing the need for frontier model CoT traces. Notably, $\pi$-Distill consistently outperforms standard SFT+RL baselines that rely on full Chain-of-Thought data. 3. Generalization: We demonstrate significant gains over baselines on $\tau$-Bench and Travel Planner, with strong generalization to eight additional out-of-distribution (OOD) on GEM tool-use tasks.
@@ 2 Background
where $K_{g}$ is the length of trajectory $g$ in tokens and $\epsilon$ is the clipping parameter.
@@ 3 Methods
$\pi$-Distill learns from teacher-generated traces by jointly improving both the teacher and the student. In contrast, on-policy self-distillation (OPSD)  [^1] We adopt the naming ”self-distillation” following concurrent work shenfeld2026selfdistillationenablescontinuallearning; zhao2026selfdistilledreasoneronpolicyselfdistillation that propose the same objective. samples trajectories from the student policy and uses a reverse KL divergence between the student and the PI-conditioned teacher as a training penalty.
@@ 3.1 Privileged Information distillation ($\pi$-Distill )
$J_{\text{Student}}(\theta)=\mathbb{E}_{\begin{subarray}{c}\mathbf{o}\sim{\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})}\\ \mathbf{s}\sim P\end{subarray}}\Big[\dfrac{{\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}(\mathbf{o}|s)}}{\operatorname{sg}\!\left({\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})}\right)}R(\mathbf{o},\mathbf{s})\Big]-\beta D_{\mathrm{KL}}\!\left(\operatorname{sg}\!\left({\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})}\right)\,\|\,{\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}(\mathbf{o}|s)}\right)$ (3)
@@ 3.2 On-Policy Self Distillation
$\pi$-Distill can be viewed as off-policy learning, where the student is trained on trajectories generated by the PI-conditioned teacher. A complementary line of work studies on-policy distillation, in which the student acts as the sampling policy and knowledge is transferred by minimizing the reverse KL between the student and teacher (agarwal2024onpolicydistillationlanguagemodels). Prior work typically utilizes a larger model as the teacher. We introduce this objective in our PI setting by instantiating it with the same shared-parameter model, where the teacher is additionally conditioned on PI. We refer to this objective as On-Policy Self-Distillation (OPSD) to maintain consistency with concurrent work (shenfeld2026selfdistillationenablescontinuallearning; zhao2026selfdistilledreasoneronpolicyselfdistillation; hübotter2026reinforcementlearningselfdistillation).
@@
Note that the updates are on-policy as the expectation is taken over $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$}. Where the reverse KL acts as a dense per-token reward measuring how closely the student matches the teacher. The full algorithm is given in algorithm˜2. We analyze OPSD in greater depth in App B and characterize the specific target distribution that the algorithm implicitly fits. Concurrent work also propose this objective, demonstrating its effectiveness in settings such as having access to ground truth answers (zhao2026selfdistilledreasoneronpolicyselfdistillation; hübotter2026reinforcementlearningselfdistillation) and continual learning (shenfeld2026selfdistillationenablescontinuallearning). In this work, we introduce and evaluate OPSD for PI transfer without ground truth, and identify the settings in which it fails.
@@ 4.1 Benchmarks
To evaluate our approach, we first employ $\tau$-Bench (yao2024taubenchbenchmarktoolagentuserinteraction), which simulates customer service interactions where agents book flights (airline domain) or assist shoppers (retail domain) by calling tools and gathering user information. To reduce computational costs, we substitute the GPT-4o user simulator with Qwen-14B. The most substantial change is that we remove the transfer_to_human_agents tool, as it consistently led to reward hacking. The resulting dataset consists of 500 training tasks restricted to the retail domain; we evaluate on 115 held-out retail tasks and 50 airline tasks, utilizing the latter to test Out-Of-Domain (OOD) generalization. Next, we utilize Travel Planner (xie2024travelplannerbenchmarkrealworldplanning), a benchmark focusing on tool use for planning. While the original repository [^2] https://github.com/OSU-NLP-Group/TravelPlanner employs a rubric-based evaluation that prioritizes "easy" constraints before checking "hard" ones, we found this setup causes policies to collapse onto undesired behaviors. To address this, we decouple the rewards so that easy constraints are tied directly to their corresponding hard constraints, for example, verifying dietary restrictions immediately after booking a restaurant. We train on the 45 training tasks and report results on the 180 publicly available held-out tasks. Finally, to probe whether training on these environments enhances tool usage to scopes far beyond the training domains, we evaluate on the GEM QA multi-turn tool-usage environment suite (liu2025gemgymagenticllms). This suite equips agents with a search tool consisting of seven environments (2Wiki (xanh2020_2wikimultihop), PopQA (mallen2023llm_memorization), TriviaQA (joshi2017triviaqalargescaledistantly), HotpotQA (hotpotqa), Bamboogle (bamboogle), NaturalQuestions (naturalquestions), and Musique (trivedi2022musiquemultihopquestionssinglehop)).
@@ 4.2 Sources of Privileged Information
To show the benefits of leveraging train-time PI, we aim to distill a frontier model using only its raw output trajectories. For this, we mine trajectories from DeepSeek-chat-v3.1 as it is open-source and allows access to its reasoning tokens. Having access to these tokens enables us to benchmark against standard settings that assume access to CoT. Enabling us to analyze the performance impact when one lacks access to the expert’s reasoning traces.
Figure 3: Deriving PI from frontier model traces. The left panel illustrates sampling trajectories from a closed-source frontier model, where full CoT reasoning is typically occluded openai2024openaio1card. We transform these raw traces into three types of PI with varying information density: (1) Tool Calls & Arguments, retaining the exact actions invoked by the frontier model; (2) Tool Calls Only, where arguments are stripped; and (3) Self-Generated Hints, where the student model summarizes the frontier trajectory into a concise hint.
@@
    Tool calls & arguments. The complete sequence of actions executed by the frontier model including function names and input arguments, e.g. GetUserDetails(Name:"Kevin Lau").
@@
    Self-generated hints. We prompt the trained model model to summarize a successful trajectory produced by the frontier model. This process can help filter inefficient actions and suggest ways to improve upon the expert’s behavior (see App. G.3 for an example).
Leveraging these three formats enables us to identify what properties allow successful transfer between teacher and student. We analyze this in section˜7 finding that the optimal configuration depends on the value of $\alpha$ and the specific attributes of the PI (e.g. KL between teacher and student).
Using this setup, we collect 15,885 successful traces for $\tau$-Bench retail and 1,986 for Travel Planner, evenly sampled across tasks. From these traces, for each training task we select the successful trajectory with the least number of steps to build the PI. In total we obtain PI for all $45$ tasks in Travel Planner and $300/500$ tasks in $\tau$-Bench. We note we exclude the self-generated hints for R1-Deepseek-Distill-LLama-8B as it consistently returns the raw input trace or tool calls as the hint.
@@ 4.3 Models and Baselines
We instantiate $\pi$-Distill with $\alpha\in\{0,0.5,1\}$ and OPSD. We compare against (i) standard RL, (ii) SFT on expert trajectories with and without CoT, and (iii) SFT followed by RL. We utilize the GRPO objective outlined in equation˜1 for all RL-based algorithms. Following vattikonda2025trainllmwebagent, we sweep over multiple SFT checkpoints for SFT+RL baselines and report results using the checkpoint that yields the strongest final performance. We use the full set of collected successful traces for all SFT baselines, as this maximized performance. For $\pi$-Distill and OPSD in $\tau$-Bench, we utilize PI whenever available, otherwise, we sample the traces with the student and perform regular RL for that goal.
@@
We run all experiments using 2 H100 GPUs with a context limit of 2525k tokens (with one exception, see App. E). We found initially that traces often exceeded this limit, thus adopting a length penalty reward that penalizes trajectories exceeding 1515k tokens (full details in App. E.1). Additionally, we observed during training that a few tokens referencing the PI consistently exhibited very high KL (e.g., the token “hint”). We thus incorporate a penalty on the frequency of these tokens. Our final experiments incorporate this penalty, though we found in practice it makes little difference for final performance (see App. D for full details). All PI is added to the task description in the system prompt, using the prompts in App. G.4. We thoroughly sweep over relevant HPs for all baselines, for a full list see App. E.2. Using the best performing set, we run three seeds of this set and report their average. In experiments, we score models using the best running average over three subsequent checkpoints, capturing both performance and stability.

### v3

@@ Abstract
Training-time privileged information (PI) can enable language models to succeed on tasks they would otherwise fail, making it a powerful tool for Reinforcement Learning (RL) in hard, long-horizon settings. However, transferring capabilities learned with PI to policies that must act without it at inference time remains a fundamental challenge. We study this problem in the context of distilling frontier models for multi-turn agentic environments, which typically hide their internal reasoning and expose only action trajectories. This breaks standard distillation pipelines, since successful behavior is observable, but the reasoning process is not. For this, we introduce $\pi$-Distill, a joint teacher-student objective that trains a PI-conditioned teacher and an unconditioned student simultaneously using the same model. Additionally, we also introduce On-Policy Self-Distillation (OPSD), an alternative approach that trains using RL with a reverse KL-penalty between the student and the PI-conditioned teacher. We show that both of these algorithms effectively distill frontier agents using action-only PI. Specifically, we find that $\pi$-Distill and, in some cases, OPSD, outperform industry standard practices (Supervised finetuning followed by RL) that assume access to full Chain-of-Thought supervision across multiple agentic benchmarks, models, and forms of PI. We complement our results with extensive analysis that characterizes the factors enabling effective learning with PI, focusing primarily on $\pi$-Distill and characterizing when OPSD is competitive.
[…]
Figure 1: Overview of the $\pi$-Distill framework. (1) Successful trajectories (not shown) are collected from a frontier agent that exposes only actions while hiding its Chain-of-Thought. (2) These trajectories are transformed into training-time privileged information (PI) and used to sample using a PI-conditioned teacher policy $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})$}. (3) The PI-conditioned teacher and an unconditioned student $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}(\mathbf{o}|s)$} share parameters and are trained jointly, enabling transfer of privileged knowledge to a test-time policy that acts without PI.
@@ 1 Introduction
Training-time PI can be particularly useful for Reinforcement Learning (RL) with LMs, where learning is contingent on the model’s ability to first succeed on a task (yue2025doesreinforcementlearningreally). That is, the model can leverage train-time PI to succeed at tasks it would otherwise fail, effectively enabling it to bootstrap its learning from these successful experiences. As noted above, the underlying challenge is transfer: training a model with PI and obtaining a test-time policy that retains those enhanced capabilities without PI.
@@
We ground our work in the task of distilling frontier models for complex multi-turn agentic settings. Typically, the industry standard for these tasks involves Supervised Fine-Tuning (SFT) on frontier model outputs followed by Reinforcement Learning (RL). Unfortunately, some model providers restrict important information, most notably the model’s full Chain-of-Thought (CoT) reasoning traces (openai2024openaio1card), providing only a summary alongside the action they intend to take. This opacity undermines standard distillation methods, as we can observe what successful agents do but not how they reason about it.
@@
Contributions 1. Algorithms for Privileged Information: We introduce $\pi$-Distill and On-Policy Self-Distillation (OPSD), two methods for training policies using train-time privileged information. 2. Distilling without CoT: We show that our methods effectively distill frontier models using actions alone, bypassing the need for frontier model CoT traces. Notably, $\pi$-Distill consistently outperforms standard SFT+RL baselines that rely on full Chain-of-Thought data. 3. Generalization: We demonstrate significant gains over baselines on $\tau$-Bench and Travel Planner, with strong generalization to eight additional out-of-distribution (OOD) on GEM tool-use tasks.
@@ 2 Background
where $K_{g}$ is the length of trajectory $g$ in tokens and $\epsilon$ is the clipping parameter. While the original GRPO objective typically employs a KL-penalty with respect to the base model, we drop this term as recent work shows it can hinder performance (shah2026comedyestimatorsklregularization).
@@ 3 Methods
Privileged Information Distillation ($\pi$-Distill) learns from teacher-generated traces by jointly improving both the teacher and the student. In contrast, on-policy self-distillation (OPSD)  [^1] We adopt the naming ”self-distillation” following concurrent work (shenfeld2026selfdistillationenablescontinuallearning; zhao2026selfdistilledreasoneronpolicyselfdistillation; hubotter2026reinforcementlearningselfdistillation) that propose the same objective. samples trajectories from the student policy and uses a reverse KL divergence between the student and the PI-conditioned teacher as a training penalty. Both algorithms are derived using a variational perspective, which we outline in detail in App. A and App. B.
@@ 3.1 Privileged Information Distillation ($\pi$-Distill)
$J_{\text{Student}}(\theta)=\mathbb{E}_{\begin{subarray}{c}\mathbf{o}\sim{\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})}\\ \mathbf{s}\sim P\end{subarray}}\Big[\dfrac{{\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}(\mathbf{o}|s)}}{\operatorname{sg}\!\left({\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})}\right)}R(\mathbf{o},\mathbf{s})\Big]-\beta D_{\mathrm{KL}}\!\left(\operatorname{sg}\!\left({\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})}\right)\,\|\,{\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}(\mathbf{o}|s)}\right).$ (3)
@@ 3.2 On-Policy Self Distillation
$\pi$-Distill can be viewed as off-policy learning, where the student is trained on trajectories generated by the PI-conditioned teacher. A complementary line of work studies on-policy distillation, in which the student acts as the sampling policy and knowledge is transferred by minimizing the reverse KL between the student and teacher (agarwal2024onpolicydistillationlanguagemodels; yang2026learningteachergeneralizedonpolicy). Prior work typically utilizes a larger model as the teacher. We introduce this objective in our PI setting by instantiating it with the same shared-parameter model, where the teacher is additionally conditioned on PI. We refer to this objective as On-Policy Self-Distillation (OPSD).
@@
Note that the updates are on-policy as the expectation is taken over $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$}. Where the reverse KL acts as a dense per-token reward measuring how closely the student matches the teacher. The full algorithm is given in algorithm˜2. We analyze OPSD in greater depth in App B and characterize the specific target distribution that the algorithm implicitly fits. Concurrent work also propose this objective, demonstrating its effectiveness in settings such as having access to ground truth answers (zhao2026selfdistilledreasoneronpolicyselfdistillation), conditioning on reflective self-feedback (hubotter2026reinforcementlearningselfdistillation), and continual learning (shenfeld2026selfdistillationenablescontinuallearning). In this work, we introduce and evaluate OPSD for PI transfer without ground truth and characterize it across a variety of PI types, identifying settings in which it fails.
@@ 4.1 Benchmarks
To evaluate our approach, we first employ $\tau$-Bench (yao2024taubenchbenchmarktoolagentuserinteraction), which simulates customer service interactions where agents book flights (airline domain) or assist shoppers (retail domain) by calling tools and gathering user information. To reduce computational costs, we substitute the GPT-4o user simulator with Qwen-14B. The most substantial change is that we remove the transfer_to_human_agents tool, as it consistently led to reward hacking. The resulting dataset consists of 500 training tasks restricted to the retail domain; we evaluate on 115 held-out retail tasks and 50 airline tasks, utilizing the latter to test Out-Of-Domain (OOD) generalization. Next, we utilize Travel Planner (xie2024travelplannerbenchmarkrealworldplanning), a benchmark focusing on tool use for planning. While the original repository [^2] https://github.com/OSU-NLP-Group/TravelPlanner employs a rubric-based evaluation that prioritizes "easy" constraints before checking "hard" ones, we found this setup causes policies to collapse onto undesired behaviors (see App. G). To address this, we decouple the rewards so that easy constraints are tied directly to their corresponding hard constraints, for example, verifying dietary restrictions immediately after booking a restaurant. We train on the 45 training tasks and report results on the 180 publicly available held-out tasks. Finally, to probe whether training on these environments enhances tool usage to scopes far beyond the training domains, we evaluate on the GEM QA multi-turn tool-usage environment suite (liu2025gemgymagenticllms). This suite equips agents with a search tool consisting of seven environments (2Wiki (xanh2020_2wikimultihop), PopQA (mallen2023llm_memorization), TriviaQA (joshi2017triviaqalargescaledistantly), HotpotQA (hotpotqa), Bamboogle (bamboogle), NaturalQuestions (naturalquestions), and Musique (trivedi2022musiquemultihopquestionssinglehop)).
@@ 4.2 Sources of Privileged Information
To show the benefits of leveraging train-time PI, we aim to distill a frontier model using only its raw output trajectories. For this, we mine trajectories from DeepSeek-chat-v3.1 as it is open-source and allows access to its reasoning tokens. Having access to these tokens allows us to benchmark against standard settings that assume access to CoT. This is an important detail, as baselines that assume access to this information should act as a soft-upper bound.
Figure 3: Deriving PI from frontier model traces. The left panel illustrates sampling trajectories from a frontier model, where full CoT reasoning is occluded (openai2024openaio1card). We transform these raw traces into three types of PI with varying information density: (1) Tool Calls & Arguments, retaining the exact actions invoked by the frontier model; (2) Tool Calls Only, where arguments are stripped; and (3) Self-Generated Hints, where the student model summarizes the frontier trajectory into a concise hint.
@@
    Tool calls & arguments. The complete sequence of actions executed by the frontier model, including function names and input arguments, e.g. GetUserDetails(Name:"Kevin Lau").
@@
    Self-generated hints. We prompt the trained model to summarize a successful trajectory produced by the frontier model. This process can help filter inefficient actions and suggest ways to improve upon the expert’s behavior (see App. H.3 for an example).
Leveraging these three formats enables us to identify what properties allow successful transfer between teacher and student. We analyze this in section˜7, finding that the optimal configuration depends on the value of $\alpha$ and the specific attributes of the PI (e.g. KL between teacher and student).
Using this setup, we collect 15,885 successful traces for $\tau$-Bench retail and 1,986 for Travel Planner, evenly sampled across tasks. From these traces, for each training task, we select the successful trajectory with the least number of steps to build the PI. In total, we obtain PI for all $45$ tasks in Travel Planner and $300/500$ tasks in $\tau$-Bench. We note we exclude the self-generated hints for R1-Deepseek-Distill-LLama-8B as it consistently returns the raw input trace or tool calls as the hint.
@@ 4.3 Models and Baselines
We instantiate $\pi$-Distill with $\alpha\in\{0,0.5,1\}$ and OPSD. We compare against (i) standard RL, (ii) SFT on expert trajectories with and without CoT, and (iii) SFT followed by RL. We utilize the GRPO objective outlined in equation˜1 for all RL-based algorithms. Following vattikonda2025trainllmwebagent, we sweep over multiple SFT checkpoints for SFT+RL baselines and report results using the checkpoint that yields the strongest final performance. We use the full set of collected successful traces for all SFT baselines, as this maximized performance. For $\pi$-Distill and OPSD in $\tau$-Bench, we utilize PI whenever available otherwise, we sample the traces with the student and perform regular RL for that goal.
@@
We run all experiments using 2 H100 GPUs with a context limit of 2525k tokens (with one exception, see App. E). We found initially that traces often exceeded this limit, thus adopting a length penalty reward that penalizes trajectories exceeding 1515k tokens (full details in App. E.2). Additionally, we observed during training that a few tokens referencing the PI consistently exhibited very high KL (e.g., the token “hint”). We thus incorporate a penalty on the frequency of these tokens. Our final experiments incorporate this penalty, though we found in practice it makes little difference for final performance (see App. D for full details). All PI is added to the task description in the system prompt, using the prompts in App. H.4. We thoroughly sweep over relevant HPs for all baselines, for a full list see App. E.3. Using the best performing set, we run three seeds of this set and report their average. In experiments, we score models using the best running average over three subsequent checkpoints, capturing both performance and stability.